# Adaptive Contrastive Decoding in Retrieval-Augmented Generation for Handling Noisy Contexts

### v1

@@ 4 Conclusion
In this work, we mainly tackle handling noisy contexts in open-domain QA on the RAG framework. This method dynamically adjusts contextual influence during decoding by quantifying uncertainty dropped by the retrieved context. Our results show that ACD improves performances across various dimensions by considering LLM’s parametric knowledge and context noisiness. These findings highlight ACD’s potential to enhance the reliability of retrieval-augmented LLMs.
@@
Appendix A Related Works
Context-augmented Generation
With the objective of closely adhering to the context provided, context-augmented generation strategies have been developed to better utilize the context and produce more factual and context-relevant responses at the inference time (Zhou et al., 2023; He et al., 2024). Especially, approaches integrated with contrastive decoding (Li et al., 2023; Malkin et al., 2022; Liu et al., 2021) yield significant enhancement by amplifying the influence of the given context at decoding step (Shi et al., 2023; Zhao et al., 2024). These context-aware contrastive decoding methods to generate responses faithful to the given context show effective performance in summarization (See et al., 2017; Narayan et al., 2018), knowledge flipping (Longpre et al., 2022), and question answering with gold-standard contexts.
Robustness in RAG Framework
Appendix B Implementation Details
B.1 Instructions
@@
B.2 Datasets
B.3 Baselines
CAD introduces a context-aware contrastive decoding approach that employs a contrastive output distribution to accentuate discrepancies in model predictions with and without context. This method effectively overrides model priors conflicting with provided context, offering significant performance enhancements in tasks requiring resolution of knowledge conflicts. MICD further enhances context grounded generation by integrating contrastive decoding with adversarial irrelevant passages.
From a computational time perspective, MICD requires three times more than conventional greedy decoding, while CAD and ACD require twice as much.
MICD proposes two usage directions, referred to as MICDF and MICDD, which offer fixed and dynamic $\alpha$, respectively. In $\text{MICD}_{D}$, $\alpha$ is assigned as the maximum token probability with context ($maxP_{wc}$) if $maxP_{wc}$ exceeds the maximum token probability without context ($maxP_{woc}$); otherwise, it is calculated as $1-maxP_{woc}$.
Throughout the experiments, fixed value of $\alpha$ is set to the value used in Zhao et al. (2024), 0.5 and 1.0 for CAD and $\text{MICD}_{F}$, respectively.
B.4 Retriever Performance
@@
Appendix C Main Results on Llama3 and Mistral
Main results on Llama3 8B and Mistral 7B are reported in Table 7.
Appendix D Full Results on Known-noisy and Unknown-gold
@@
Appendix E AUROC between $\alpha$ and Context Noisiness
@@
Appendix F Case Study on $\alpha_{\text{ACD}}$
@@
In case of Known-noisy, when the model generates the answer correctly even without the given context, the retrieved noisy context yields relatively higher entropy, resulting in $\alpha_{ACD}$ value of 0.3483.
Conversely, in the case of Unknown-gold, the model’s generated answer is incorrect, aligning with a relatively high entropy value of 6.6748. In this scenario, the retrieved gold context guides the model to correctly answer the question, which is reflected in a relatively lower entropy value of 1.5628. Thus, the value of $\alpha_{ACD}$, adjusted with these entropy values, yields a relatively higher weight on the context at 0.8103.
Appendix G Knowledge Conflict
In case where the provided context is different from what the model knows but relevant to the question, we verify whether the two decoding methods with dynamic weight, ACD and $\text{MICD}_{D}$, can generate responses without considering it as a noisy context with NQ-Swap dataset (Longpre et al., 2022).
NQ-swap (Longpre et al., 2022) is constructed by replacing the answer entity span in the original gold context of NQ questions with a random entity, to verify if the knowledge is successfully flipped. We use the questions and swapped contexts provided in Hong et al. (2024) with 3650 samples excluding 5 examples used as few-shot prompts and samples with the context in a tabular format due to the limited context length. In the case of NQ-SWAP, each data point has a given context. Since it is a task that does not use a retriever, for MICD, we use the fixed negative context taken from the MICD as an adversarial context. MICD reports that the performance difference between fixed negative and the most-distant context is negligible.
Figure 3 illustrates that ACD consistently exceeds the performance of $\text{MICD}_{D}$ across all models and achieves results comparable to open-book regular decoding. The results indicate that the ACD’s approach remains effective even in settings where the context is relevant to the question but contradicts the parametric knowledge.
Datasets ($\rightarrow$) TriviaQA NQ PopQA
Methods ($\downarrow$) All $\text{Subset}_{\text{Gold}}$ $\text{Subset}_{\text{Noisy}}$ All $\text{Subset}_{\text{Gold}}$ $\text{Subset}_{\text{Noisy}}$ All $\text{Subset}_{\text{Gold}}$ $\text{Subset}_{\text{Noisy}}$
Llama3 8B
$\text{Reg}_{Cls}$ 61.67 - - 28.34 - - 32.65 - -
$\text{Reg}_{Opn}$ 61.27 86.94 36.02 33.30 63.10 14.40 39.73 82.95 8.64
CAD 49.70 72.45 27.31 29.17 58.39 10.64 35.86 76.82 6.40
MICDF 61.01 85.40 37.00 27.62 51.89 12.22 37.99 77.12 9.85
MICDD 64.01 86.08 42.28 30.72 53.96 15.98 41.35 79.32 14.04
ACD 66.32 89.20 43.81 35.48 62.03 18.65 43.25 84.48 13.60
Mistral 7B
$\text{Reg}_{Cls}$ 63.72 - - 29.64 - - 29.04 - -
$\text{Reg}_{Opn}$ 60.45 86.85 34.48 32.55 64.67 12.18 38.28 81.26 7.36
CAD 44.69 66.89 22.85 24.10 52.25 6.25 33.93 73.95 5.15
MICDF 63.33 88.43 38.62 31.80 61.10 13.22 36.58 76.00 8.23
MICDD 66.97 89.24 45.05 33.24 57.89 17.61 39.87 78.46 12.11
ACD 67.82 90.16 45.83 35.37 62.17 18.38 41.47 82.90 11.68
Table 7: EM accuracy of full data (All) and subsets with gold context ($\text{Subset}_{\text{Gold}}$) and with noisy context ($\text{Subset}_{\text{Noisy}}$).

### v2

@@ 5 Conclusion
In this work, we mainly tackle handling noisy contexts in open-domain QA on the RAG framework. Our proposed method, ACD, dynamically adjusts contextual influence during decoding by quantifying the model’s uncertainty that is either reduced or increased by the retrieved context. Our results show that ACD improves performances across various dimensions by considering the LLM’s parametric knowledge and context noisiness. These findings highlight ACD’s potential to enhance the reliability of retrieval-augmented generation.
@@
Acknowledgement
This work was partly supported by SNU-NAVER Hyperscale AI Center and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), No.RS-2020-II201373, Artificial Intelligence Graduate School Program (Hanyang University), NO.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)]
Appendix
Appendix A Implementation Details
A.1 Instructions
@@ A.2 Datasets
A.3 Baselines
Baselines using regular greedy decoding are evaluated under two different settings. In the closed-book setting, only the question is provided. In the open-book setting, the retrieved context is employed. The same top-1 retrieved context is utilized for every baseline and ACD.
CAD introduces a context-aware contrastive decoding approach that employs a contrastive output distribution to accentuate discrepancies in model predictions with and without context. This method effectively overrides model priors conflicting with provided context, offering significant performance enhancements in tasks requiring resolution of knowledge conflicts. MICD further enhances context grounded generation by integrating contrastive decoding with adversarial irrelevant passages. From a computational time perspective, MICD requires three times more than conventional greedy decoding, while CAD and ACD require twice as much.
MICD proposes two usage directions, referred to as MICDF and MICDD, which offer fixed and dynamic $\alpha$, respectively. MICDD determines $\alpha$ in use by comparing the highest token probabilities with and without given context. Throughout the experiments, fixed value of $\alpha$ is set to the value used in Zhao et al. (2024), 0.5 and 1.0 for CAD and $\text{MICD}_{F}$, respectively.
A.4 Retriever Performance
@@
A.5 Knowledge Conflict
For the NQ-swap dataset, we utilize the questions and entity-swapped contexts provided in Hong et al. (2024), which includes 3,650 samples. This total excludes 5 few-shot samples and those with contexts presented in a tabular format due to the limited context length. In the case of NQ-swap, each data point has a given context. Since it is a task that does not use a retriever, for MICD, we use the fixed negative context taken from the MICD as an adversarial context. MICD reports that the performance difference between fixed negative and the most distant context is negligible.
Appendix B Results
B.1 Results on Known-noisy and Unknown-gold
@@
B.2 AUROC between Adaptive Weight and Context Noisiness
@@
Appendix C Additional Analysis
C.1 Upper-bound of Alpha
NQ TriviaQA PopQA
Llama2-7B
$\alpha_{ACD}$ 32.91 64.85 41.29
$\alpha_{oracle}$ 35.35 (+2.44) 65.31 (+0.46) 44.10 (+2.81)
Llama2-13B
$\alpha_{ACD}$ 36.12 67.37 43.35
$\alpha_{oracle}$ 38.75 (+2.63) 68.19 (+0.82) 47.01 (+3.66)
Llama3 8B
$\alpha_{ACD}$ 35.48 66.32 43.25
$\alpha_{oracle}$ 36.98 (+1.50) 66.10 (-0.22) 46.47 (+3.22)
Mistral 7B
$\alpha_{ACD}$ 35.37 67.82 41.47
$\alpha_{oracle}$ 38.37 (+3.00) 67.29 (-0.53) 44.53 (+3.06)
Table 6: EM score comparison between ACD ($\alpha_{ACD}$) and ACD with oracle alpha value ($\alpha_{oracle}$).
In our approach, the parameter $\alpha$ is expected to be close to 1 when the retrieved context contains information that helps answer the given question, and close to 0 otherwise. To evaluate the upper-bound performance of ACD, we assume that we have prior knowledge of whether the context in use is gold or noisy. Under this assumption, we fix the $\alpha$ value to 1.0 if the context is gold and to 0.0 if the context is noisy.
For TriviaQA dataset, the performance of ACD is comparable to $\alpha_{oracle}$, with less than 1 point difference (Table 6). NQ and PopQA show a difference of approximately 2-3 points, indicating that the method for calculating the $\alpha$ weight could be further enhanced in future research.
C.2 Case Study
@@
In the case of Known-noisy, when the model generates the answer correctly even without the given context, the retrieved noisy context yields relatively higher entropy, resulting in $\alpha_{ACD}$ value of 0.3483. Conversely, in the case of Unknown-gold, the model’s generated answer is incorrect, aligning with a relatively high entropy value of 6.6748. In this scenario, the retrieved gold context guides the model to correctly answer the question, which is reflected in a relatively lower entropy value of 1.5628. Thus, the value of $\alpha_{ACD}$, adjusted with these entropy values, yields a relatively higher weight on the context at 0.8103.